$SPAD/src/input spadprof.input

The Axiom Team December 3, 2016

Abstract

Contents

2

— * —

-- This file is read on start-up by the Scratchpad system.  It is
-- most commonly used to customize the user's environment via the
-- )set system commands.  This version is for AIX.
-- read the system synonym file for system commands, if it exists
)read "synonym.input" )ifthere )quiet
-- read the user's synonym file for system commands, if it exists
)read "mysyns.input"  )ifthere )quiet
-- do not enter lisp break on error
)set break nobreak
-- start up the workspace history facility
)set history on
-- use the simple character set
)set output char plain
-- everyone uses this
)set userlevel development
-- make the highlighting characters go away
)set mess high off

———-

References^{[1] nothing}

4